# Foundations of Quantum Granular Computing with Effect-Based Granules, Algebraic Properties and Reference Architectures

### v1

@@ 6 Models and Architectures
Within this landscape, we single out three reference architectures for QGC. The first is the Measurement-Driven Granular Partitioning (MDGP) pattern, where classical granules are formed at the data level and then lifted to quantum effects. The second is Variational Effect Learning (VEL), in which quantum granules are parameterized effects learned from data. The third is a Hybrid Classical–Quantum (HCQ) pipeline that combines classical granulation with a QGDS decision layer. These architectures specify how quantum granules are defined or learned and how they interface with the QGDS layer introduced in Sec. 5.
@@ 6.1 Hybrid models: classical granulation with quantum encoding
Hybrid models leverage established techniques of classical GrC to preprocess data into granules, which are then encoded into quantum states for further manipulation. They are hybrid in the sense that the granulation and part of the reasoning remain explicitly classical (fuzzy, rough, MFL-based), while the subsequent evaluation and decision are carried out in the quantum, operator-valued domain. A typical hybrid QGC pipeline comprises three stages:
@@
    Granulation step. Given a dataset $X$, a classical granulation procedure partitions or clusters $X$ into granules $\{G_{1},\dots,G_{m}\}$, each associated with membership degrees $\mu_{i}(x)$ for $x\in X$. These memberships capture interpretable, graded inclusion of data points in each granule and can be obtained, for example, from fuzzy clustering, rough approximations, or MFL-based mediative schemes.
@@
This approach enables the direct use of quantum algorithms (e.g., qPCA, amplitude amplification, variational classifiers) on structured representations produced by classical GrC, combining the interpretability of classical granules with the potential expressiveness and parallelism of quantum state manipulations. In the broader QGDS setting, the hybrid architecture thus uses classical granulation to shape the input space and quantum effect-based granules to implement the final decision layer, making the overall model explicitly interpretable yet genuinely quantum-enhanced.
@@ 6.2 Purely quantum models
Purely quantum models define granules directly within Hilbert space, without relying on classical preprocessing or external membership functions. In this regime, granules correspond to projectors or effects $\{P_{i}\}$ or $\{E_{i}\}$ satisfying $\sum_{i}P_{i}=I$ (PVM) or $\sum_{i}E_{i}=I$ (POVM), and data are interpreted solely as quantum states $\rho$ processed through granular measurements. This paradigm aligns with quantum machine-learning approaches in which the model is fully encoded in operator-valued transformations and measurements [2, 31]. From the QGC viewpoint, the basic building block is, therefore, a measurement-driven granular model in which decisions are derived directly from the statistics of a PVM or POVM.
@@
A quantum system is probed using a PVM or POVM, where each outcome corresponds to a quantum granule $(P_{i},p_{i})$ or $(E_{i},p_{i})$:
@@
for any subset $S$ of outcomes. Since $\sum_{i}E_{i}=I$ and $0\preceq E_{i}\preceq I$, each $E_{S}$ is again an effect, mirroring multilevel approximations in classical GrC but now expressed through operator sums [10, 4]. These hierarchies provide the quantum analogue of classical refinement and coarsening of granules.
@@
This approach eliminates the need for external membership functions: degrees of belonging are intrinsically defined through Born probabilities. Moreover, entangled states naturally support joint granules acting across subsystems, enabling the representation of correlations and contextual structures that have no classical analogue. Such non-classical granularity underpins the expressive power of purely quantum models and is central to variational quantum architectures [6].
@@ 6.3 QGDS as a unifying decision layer
We now view the architectures above through the lens of Quantum Granular Decision Systems (QGDS), as introduced in Sec. 5. A QGDS provides a unifying decision layer that can be instantiated on top of both hybrid models (Sec. 6.1) and purely quantum models (Sec. 6.2).
@@
Figure 3: Hybrid pipeline for Quantum Granular Decision Systems (QGDS). Classical granulation structures the input, which is then encoded as a quantum state, processed through granular measurements (PVM/POVM), and mapped to a final decision by a Bayes-style rule.
@@ 6.4 Variational Effect Learning (VEL)
While Measurement-Driven Granular Partitioning (MDGP) fixes the granular effects a priori, Variational Effect Learning (VEL) treats them as trainable quantum parameters. In a VEL architecture, the QGDS granules are realized by a parametrized POVM $\{E_{j}(\theta)\}_{j=1}^{m}$ acting on the encoded state $\rho(x)$, and the parameters $\theta$ are optimized from data under the physical constraints $0\preceq E_{j}(\theta)\preceq I$ and $\sum_{j=1}^{m}E_{j}(\theta)=I$.
A convenient parametrization is to start from a fixed POVM $\{F_{j}\}_{j=1}^{m}$ (e.g., orthogonal projectors or block-diagonal templates) and conjugate it by a variational unitary $U(\theta)$:
@@
which fit directly into the generic QGDS decision rule
@@
where $p(x^{(n)};\theta)=(p_{1}(x^{(n)};\theta),\dots,p_{m}(x^{(n)};\theta))$ and $\ell$ is a suitable loss function (e.g., cross-entropy for classification or a margin-based loss). The optimization is carried out on a classical processor using gradient-based or gradient-free methods, while all constraints on $\{E_{j}(\theta)\}$ are enforced through the parametrization in (41).
Conceptually, VEL provides a data-driven mechanism to learn quantum granules that remain physically valid effects while adapting to the structure of the task. In the commutative (classical) limit, where all $\rho(x)$ and $E_{j}(\theta)$ share a common eigenbasis, VEL reduces to learning classical soft granules over a Boolean island (Theorem 2). In genuinely non-commutative settings, VEL can exploit Hilbert-space geometry and entanglement to realize granular decision boundaries with no direct classical analogue, while remaining compatible with NISQ hardware via shallow, structured ansätze for $U(\theta)$ [6].
@@ 7 Case Studies
This section presents compact case studies illustrating how quantum granules behave in simple quantum systems. These examples connect the abstract definitions of effects and POVMs with concrete qubit-level behavior and with hybrid classical–quantum pipelines. We focus on one- and two-qubit systems because they provide clear geometric and algebraic intuition.
@@ 7.1 Granulation of pure qubit states
If $E$ is projective, the granular region becomes a crisp spherical cap on the Bloch sphere. If $E$ is non-projective, the granular region becomes “fuzzy”, with $0<p_{\rho}(E)<1$ governed by the eigenvalues of $E$.
Varying the effect across a family $\{E_{j}\}$ induces a partition of the Bloch sphere in the QGC sense, with each $E_{j}$ acting as a quantum granule over the space of pure qubit states.
@@ 7.2 Granulation of mixed qubit states
Mixed qubit states interpolate between pure states and the maximally mixed state $I/2$. A standard parameterization is
@@
where $\vec{r}$ is the Bloch vector in the usual Bloch-sphere representation [25].
Any qubit effect has a Bloch representation
@@
where $\vec{e}=(e_{x},e_{y},e_{z})$, $\vec{\sigma}=(X,Y,Z)$, and $\|\vec{e}\|=\sqrt{e_{x}^{2}+e_{y}^{2}+e_{z}^{2}}$ are the Euclidean norms of $\vec{e}$. For $E$ to be a valid effect (i.e., $0\preceq E\preceq I$), the Bloch parameters must satisfy
@@ 7.3 Two-qubit granulation: parity effects
Two-qubit systems permit simple illustrations of commuting and non-commuting granules. Define the parity effects:
@@
These operators satisfy $E_{\mathrm{even}}+E_{\mathrm{odd}}=I$ and commute. In fact, $E_{\mathrm{even}}$ and $E_{\mathrm{odd}}$ are projectors, so they form a PVM and represent sharp, complementary granules corresponding to even and odd parity.
@@
Thus, granulation reduces to classical parity classification on a Boolean island (Theorem 2). This abstraction is useful in error detection and syndrome extraction, where only the global parity granule is relevant.
@@ 7.5 Hybrid classical–quantum example
In the QGDS framework, this pipeline follows the stages already depicted schematically in Fig. 3: classical data are granulated, encoded as quantum states, processed by quantum effects, and finally aggregated into a decision by a classical rule.
This pipeline shows how QGC integrates classical interpretability with quantum operator structure, enabling granular reasoning under incompatibility and contextuality while keeping the decision stage compatible with existing granular and fuzzy methodologies.
@@ 8 Discussion
The development of Quantum Granular Computing (QGC) provides an operator-theoretic extension of classical granulation that accommodates incompatibility, contextuality, and measurement-induced disturbance. This section discusses the interpretability of quantum granules, their relationship to classical granular models, and the implications of working in a non-distributive algebraic setting. We also outline practical considerations for near-term implementations.
@@ 8.1 Interpretability of quantum granules
    their eigenstructure corresponds to meaningful directions (e.g., Bloch-sphere axes or physically interpretable observables),
@@
    they arise from a physically motivated PVM/POVM (e.g., energy levels, parity, or syndrome measurements),
@@ 8.2 Relation to classical granulation models
    MFL and mediative reasoning. Mediative interactions in Mediative Fuzzy Logic correspond to compatibility and conflict between granules; QGC extends this idea to operator-mediated interactions (via non-commuting effects and channels) that have no classical counterpart [23, 5].
Quantum granules reduce to classical granules whenever the family of effects commutes. This phenomenon—formalized by the Boolean-islands theorem—provides a clear bridge between quantum and classical granular computation: on commuting islands, QGC collapses to a classical probability space; outside them, genuinely quantum features such as contextuality and entanglement emerge [3, 30].
@@ 8.3 Non-distributive effects and contextuality
    state disturbance induced by granule evaluation, which changes subsequent membership degrees in a controlled way (e.g., Lüders refinement).
@@ 8.5 Limitations and practical considerations
    Hardware restrictions. Current devices constrain the depth and connectivity of ansätze used in VEL and QGDS. This limits the complexity of learnable granules and may favor shallow, structured architectures over generic parameterizations, in line with broader observations on variational quantum algorithms [6].
@@
    Interpretability of learned granules. While structured ansätze improve readability, arbitrary learned effects may lack intuitive meaning. Additional constraints (sparsity, block structure, or commutation with physically meaningful observables) may be needed to maintain semantic transparency, particularly in decision-support applications where explainability is critical.
@@ 9 Conclusions
This paper has developed the foundations of Quantum Granular Computing (QGC), extending classical granular computing—including fuzzy, rough, and shadowed granules—to the quantum regime. The key step is to model quantum granules as effects on a finite-dimensional Hilbert space, so that granular memberships are given by Born probabilities. This operator-theoretic viewpoint provides a common language for sharp (projective) and soft (non-projective) granules and embeds granulation directly into the standard formalism of quantum information theory.
@@
On this basis, we established several foundational results for effect-based quantum granules. We characterized basic normalization and monotonicity properties, identified Boolean islands of commuting granules where classical probabilistic reasoning is recovered, and analyzed granular refinement under Lüders updates. We also described the evolution of granules under quantum channels via the Heisenberg adjoint, showing how noisy dynamics “dress” ideal granules into effective ones. These results clarify how granular structure behaves in non-commutative settings and how it is affected by measurement and open-system dynamics.
@@
We further connected QGC with quantum detection and estimation theory by interpreting optimal binary decisions as optimal decision granules of Helstrom type. In this formulation, the Helstrom projector becomes a canonical quantum granule that implements Bayes-optimal soft decisions, providing a direct analogue of fuzzy-style classifiers within an operator-theoretic framework. This link illustrates how granular reasoning and optimal quantum decision rules can be treated within a single formalism.
@@
Building on these theoretical foundations, we introduced Quantum Granular Decision Systems (QGDS) as an operational model for decision-making in QGC. We outlined three reference architectures: Measurement-Driven Granular Partitioning (MDGP), Variational Effect Learning (VEL), and Hybrid Classical–Quantum pipelines (HCQ). Together, these architectures describe how quantum granules can be specified, learned, and integrated with classical components in a structured way, while remaining compatible with near-term quantum hardware and standard quantum-information primitives.
@@
Compact case studies on qubit granulation, two-qubit parity effects, and Helstrom-style soft decisions illustrated how the proposed framework behaves in simple scenarios. These examples showed that QGC can reproduce fuzzy-like behavior (soft granules, graded memberships, and smooth decision boundaries) while exploiting non-commutativity, contextuality, and entanglement. In this sense, QGC does not simply mimic classical granulation in Hilbert space, but enlarges the design space of granular models by allowing genuinely quantum phenomena.
@@
Overall, the contribution of this work is a unified and mathematically grounded foundation for QGC that connects classical granular models, operator theory, and modern quantum information science. In particular, our operator-based formulation provides a unified treatment of sharp and soft quantum granules while recovering fuzzy, rough, and shadowed granules as commutative and projective special cases of quantum granules, as formalized by our Boolean-islands theorem. This makes classical GrC a genuine subtheory of QGC, rather than a loosely related analogy, and provides a basis for future research on quantum information processing, granular reasoning, and intelligent systems, particularly in domains where contextuality and non-commutativity are intrinsic rather than incidental features of the data and the decision process.

### v2

@@ 6 Models and Architectures
Within this framework, we focus on three reference architectures for QGC. The first is the Measurement-Driven Granular Partitioning (MDGP) pattern, where classical granules are formed at the data level and subsequently lifted to quantum effects. The second is Variational Effect Learning (VEL), in which quantum granules are parameterized effects learned from data. The third is a Hybrid Classical–Quantum (HCQ) pipeline that combines classical granulation with a QGDS decision layer. These architectures specify how quantum granules are defined or learned and how they interface with the QGDS layer introduced in Sec. 5.
@@ 6.1 Hybrid models: classical granulation with quantum encoding
Hybrid models use established techniques of classical GrC to preprocess data into granules, which are then encoded into quantum states for further manipulation. They are hybrid in the sense that granulation and part of the reasoning remain explicitly classical (fuzzy, rough, shadowed, MFL-based, or other GrC formalisms), while the subsequent evaluation and decision are carried out in the quantum, operator-valued domain.
A typical hybrid QGC pipeline comprises three stages:
@@
    Granulation step. Given a dataset $X$, a classical granulation procedure partitions or clusters $X$ into granules $\{G_{1},\dots,G_{m}\}$, each associated with membership degrees $\mu_{i}(x)$ for $x\in X$. These memberships capture interpretable, graded inclusion of data points in each granule and can be obtained, for example, from fuzzy clustering, rough approximations, shadowed sets, MFL-based mediative schemes, or related GrC constructions.
@@
This architecture allows the direct use of quantum algorithms (for instance, quantum principal component analysis, amplitude amplification, or variational classifiers) on structured representations produced by classical GrC, combining the interpretability of classical granules with the potential expressiveness and parallelism of quantum state manipulations. In the broader QGDS setting, the hybrid architecture uses classical granulation to shape the input space and quantum effect-based granules to implement the final decision layer, so that the overall model remains explicitly interpretable while being genuinely quantum-enhanced.
@@ 6.2 Purely quantum models
Purely quantum models define granules directly within Hilbert space, without relying on classical preprocessing or external membership functions. In this regime, granules correspond to projectors or, more generally, to effects $\{P_{i}\}$ or $\{E_{i}\}$ satisfying $\sum_{i}P_{i}=I$ (PVM) or $\sum_{i}E_{i}=I$ (POVM), with $0\preceq E_{i}\preceq I$, and data are interpreted solely as quantum states $\rho$ processed through granular measurements. This paradigm aligns with quantum machine-learning approaches in which the model is fully encoded in operator-valued transformations and measurements [3, 31]. From the QGC viewpoint, the basic building block is therefore a measurement-driven granular model in which decisions are derived directly from the statistics of a PVM or POVM.
@@
A quantum system is probed using a PVM or POVM, where each outcome corresponds to a quantum granule specified by $(P_{i},p_{i})$ or $(E_{i},p_{i})$:
@@
for any subset $S$ of outcomes. Since $\sum_{i}E_{i}=I$ and $0\preceq E_{i}\preceq I$, each $E_{S}$ is again an effect, which mirrors multilevel approximations in classical GrC but now expressed through operator sums [11, 5]. These hierarchies provide the quantum analogue of classical refinement and coarsening of granules.
@@
This approach eliminates the need for external membership functions: degrees of belonging are intrinsically defined through Born probabilities. Moreover, entangled states naturally support joint granules acting across subsystems, which enables the representation of correlations and contextual structures with no classical analogue. Such non-classical granularity underpins the expressive power of purely quantum models and is central to variational quantum architectures [7].
@@ 6.3 QGDS as a unifying decision layer
The architectures above can be viewed through the lens of Quantum Granular Decision Systems (QGDS), as introduced in Sec. 5. A QGDS provides a unifying decision layer that can be instantiated on top of both hybrid models (Sec. 6.1) and purely quantum models (Sec. 6.2).
@@
Figure 3: Hybrid pipeline for a Quantum Granular Decision System (QGDS). Classical granulation structures the input, which is then encoded as a quantum state, processed through granular measurements (PVM/POVM), and mapped to a final decision by a Bayes-type rule acting on the resulting membership vector.
@@ 6.4 Variational Effect Learning (VEL)
Measurement-Driven Granular Partitioning (MDGP) fixes the granular effects a priori, whereas Variational Effect Learning (VEL) treats them as trainable quantum parameters. In a VEL architecture, the QGDS granules are realized by a parametrized POVM $\{E_{j}(\theta)\}_{j=1}^{m}$ acting on the encoded state $\rho(x)$, and the parameters $\theta$ are optimized from data under the physical constraints
$0\preceq E_{j}(\theta)\preceq I,\qquad\sum_{j=1}^{m}E_{j}(\theta)=I.$ (47)
A convenient parametrization is to start from a fixed POVM $\{F_{j}\}_{j=1}^{m}$ (for example, orthogonal projectors or block-diagonal templates) and conjugate it by a variational unitary $U(\theta)$:
@@
which enter directly into the generic QGDS decision rule
@@
where $p(x^{(n)};\theta)=(p_{1}(x^{(n)};\theta),\dots,p_{m}(x^{(n)};\theta))$ and $\ell$ is a suitable loss function (for example, cross-entropy for classification or a margin-based loss). The optimization is carried out on a classical processor using gradient-based or gradient-free methods, while all constraints on $\{E_{j}(\theta)\}$ are enforced through the parametrization in (48).
VEL thus provides a data-driven mechanism to learn quantum granules that remain physically valid effects and adapt to the structure of the task. In the commutative (classical) limit, where all $\rho(x)$ and $E_{j}(\theta)$ share a common eigenbasis, VEL reduces to learning classical soft granules over a Boolean island (Theorem 2). In genuinely non-commutative settings, VEL can exploit Hilbert-space geometry and entanglement to realize granular decision boundaries with no direct classical analogue, while remaining compatible with NISQ hardware via shallow, structured ansätze for $U(\theta)$ [7].
@@ 7 Case Studies
This section presents concise case studies illustrating how quantum granules behave in simple quantum systems. These examples connect the abstract definitions of effects and POVMs with concrete qubit-level behavior and with hybrid classical–quantum pipelines. We focus on one- and two-qubit systems because they provide clear geometric and algebraic intuition.
@@ 7.1 Granulation of pure qubit states
If $E$ is projective, the granular region corresponds to a crisp spherical cap on the Bloch sphere. If $E$ is non-projective, the granular region is fuzzy, with $0<p_{\rho}(E)<1$ governed by the eigenvalues of $E$.
When the effects form a POVM $\{E_{j}\}$, the family of quantum granules $\{E_{j}\}$ induces a soft partition of the Bloch sphere in the QGC sense, with each $E_{j}$ acting as a quantum granule over the space of pure qubit states.
@@ 7.2 Granulation of mixed qubit states
Mixed qubit states interpolate between pure states and the maximally mixed state $I/2$. A standard parametrization is
@@
where $\vec{r}=(r_{x},r_{y},r_{z})$ is the Bloch vector in the usual Bloch-sphere representation [24].
Any qubit effect can be written in Bloch form as
@@
where $\alpha\in\mathbb{R}$, $\vec{e}=(e_{x},e_{y},e_{z})\in\mathbb{R}^{3}$, $\vec{\sigma}=(X,Y,Z)$, and
$\|\vec{e}\|=\sqrt{e_{x}^{2}+e_{y}^{2}+e_{z}^{2}}$ (56)
is the Euclidean norm of $\vec{e}$. For $E$ to be a valid effect (i.e., $0\preceq E\preceq I$), the Bloch parameters must satisfy
@@ 7.3 Two-qubit granulation: parity effects
Two-qubit systems provide simple illustrations of commuting and non-commuting granules. Define the parity effects
@@
These operators satisfy $E_{\mathrm{even}}+E_{\mathrm{odd}}=I$ and commute. Moreover, since $(Z\otimes Z)^{2}=I\otimes I$, both $E_{\mathrm{even}}$ and $E_{\mathrm{odd}}$ are projectors, so they form a PVM and represent sharp, complementary granules corresponding to even and odd parity.
@@
Thus, parity granulation reduces to classical binary classification on a Boolean island (Theorem 2). This representation appears, for instance, in error detection and syndrome extraction, where only the global parity granule is relevant.
@@ 7.4 Fuzzy-like decision via optimal granules
$P_{\mathrm{succ}}(E^{\star})=\tfrac{1}{2}\bigl(1+\|\Delta\|_{1}\bigr),$ (63)
where $\|\Delta\|_{1}$ denotes the trace norm of $\Delta$.
@@ 7.5 Hybrid classical–quantum example
In the QGDS framework, this pipeline follows the stages depicted schematically in Fig. 3: classical data are granulated, encoded as quantum states, processed by quantum effects, and finally aggregated into a decision by a classical rule.
This setting illustrates how QGC integrates classical interpretability with quantum operator structure, enabling granular reasoning under incompatibility and contextuality while keeping the decision stage compatible with existing granular and fuzzy methodologies.
@@ 8 Discussion
The development of Quantum Granular Computing (QGC) provides an operator-theoretic extension of classical granulation that accommodates incompatibility, contextuality, and measurement-induced disturbance. This section discusses the interpretability of quantum granules, their relationship to classical granular models, and the implications of working in a non-distributive algebraic setting. It also outlines practical considerations for near-term implementations.
@@ 8.1 Interpretability of quantum granules
    their eigenstructure corresponds to meaningful directions (for example, Bloch-sphere axes or physically interpretable observables),
@@
    they arise from a physically motivated PVM/POVM (for example, energy levels, parity, or syndrome measurements),
@@ 8.2 Relation to classical granulation models
    MFL and mediative reasoning. Mediative interactions in Mediative Fuzzy Logic correspond to compatibility and conflict between granules. QGC extends this idea to operator-mediated interactions, via non-commuting effects and channels, that have no classical counterpart [22, 6].
Quantum granules reduce to classical granules whenever the family of effects commutes. This phenomenon, formalized by the Boolean-islands result in Theorem 2, provides a clear bridge between quantum and classical granular computation: on commuting islands, QGC collapses to a classical probability space; outside them, genuinely quantum features such as contextuality and entanglement emerge [4, 29].
@@ 8.3 Non-distributive effects and contextuality
    state disturbance induced by granule evaluation, which changes subsequent membership degrees in a controlled way (for example, through Lüders refinement).
@@ 8.5 Limitations and practical considerations
    Hardware restrictions. Current devices constrain the depth and connectivity of ansätze used in VEL and QGDS. This limits the complexity of learnable granules and may favor shallow, structured architectures over generic parametrizations, in line with observations on variational quantum algorithms [7].
@@
    Interpretability of learned granules. While structured ansätze improve readability, arbitrary learned effects may lack intuitive meaning. Additional constraints (for example, sparsity, block structure, or commutation with physically meaningful observables) may be needed to maintain semantic transparency, particularly in decision-support applications where explainability is critical.
@@ 9 Conclusions
This paper has outlined an operator-theoretic framework for Quantum Granular Computing (QGC), extending classical granular computing, including fuzzy, rough, and shadowed granules, to the quantum regime. A central modeling choice is to represent quantum granules as effects on a finite-dimensional Hilbert space, so that granular memberships are given by Born probabilities. This operator-theoretic viewpoint offers a common language for sharp (projective) and soft (non-projective) granules and embeds granulation directly into the standard formalism of quantum information theory.
@@
On this basis, we collected several basic properties of effect-based quantum granules. We characterized normalization and monotonicity properties, formulated a “Boolean islands” result for commuting families in which classical probabilistic reasoning is recovered, and analyzed granular refinement under Lüders updates. We also described the evolution of granules under quantum channels via the Heisenberg adjoint, thereby clarifying how noisy dynamics transform ideal granules into effective ones in non-commutative settings.
@@
We also highlighted a connection between QGC and quantum detection and estimation theory by interpreting optimal binary decisions as Helstrom-type decision granules. In this formulation, the Helstrom projector provides a natural example of a quantum granule that realizes Bayes-optimal soft decisions, suggesting an analogy with fuzzy-style classifiers within an operator-theoretic framework. This link illustrates how granular reasoning and optimal quantum decision rules can be treated within a single formalism.
@@
Building on these theoretical elements, we introduced Quantum Granular Decision Systems (QGDS) as an operational model for decision-making in QGC. We outlined three reference architectures: Measurement-Driven Granular Partitioning (MDGP), Variational Effect Learning (VEL), and Hybrid Classical–Quantum pipelines (HCQ). Together, these architectures illustrate possible ways in which quantum granules can be specified, learned, and integrated with classical components in a structured way, while remaining compatible with near-term quantum hardware and standard quantum-information primitives.
@@
Compact case studies on qubit granulation, two-qubit parity effects, and Helstrom-style soft decisions served to illustrate how the proposed framework behaves in simple scenarios. These examples suggest that QGC can reproduce fuzzy-like behavior (soft granules, graded memberships, and smooth decision boundaries) while also exploiting non-commutativity, contextuality, and entanglement. In this sense, QGC does not simply mimic classical granulation in Hilbert space, but appears to enlarge the design space of granular models by allowing genuinely quantum phenomena.
@@
Overall, we have proposed a mathematically grounded framework for QGC that connects classical granular models, operator theory, and modern quantum information science. In particular, the operator-based formulation provides a unified treatment of sharp and soft quantum granules while recovering fuzzy, rough, and shadowed granules as commutative and projective special cases of quantum granules, as formalized by the Boolean-islands result in Theorem 2. At the level considered here, this supports the view that core models of classical GrC can be embedded as commutative, projective instances of QGC, and it provides a basis for further work on quantum information processing, granular reasoning, and intelligent systems in domains where contextuality and non-commutativity are intrinsic features of the data and the decision process.